OpenStudy (anonymous):

Subtraction Equation: q-18-5=20

OpenStudy (anonymous):

q - 18 - 5 = 20 --> q = 20 + 18 + 5 = 43.

OpenStudy (poopsiedoodle):

or, q-23=20. 20+23=43. either one works.

OpenStudy (omnomnom):

OR well actually the first one is the easiest way to explain... so i cant give you a better explanation then that T_T

OpenStudy (anonymous):

well okay lolz

OpenStudy (omnomnom):

Lol :D